find the difference quotient of f(x) = 2x/x-1
Answer #1

Sol diffierence quotient f(x+h)-f(x) f(x)= 20 f(a+h) = 2(8th) 1-1 xth-/ 236 + 2h x+ha x-1 = = (36-1)(2x426 ) – 23c [ xth-1] (